Supply: 50mL of 15% and 85%
Order: 40%
what is the ml 85%
what is the ml 15%

Answer by stanbon(75887)

Supply: 50mL of 15% and 85%
Order: 40%
------
Equation:
0.15x + 0.85(50-x) = 0.40*50
-----
15x + 85*50 - 85x = 40*50
-70x = -45*50
x = (5/7)45
x = 32.14 mL (amt. of 15% solution needed)
50-x = 17.86 mL (amt. 85% solution needed)
=============================
Cheers,
Stan H.
-----------